# Determining Energy Efficiency Sweet Spots in Production LLM Inference

### v1

@@ 6.3. Energy Efficiency Evaluation
In Figure 2 we show the Energy Efficiency $\frac{\text{Tokens}}{\text{Joule}}]$ for each LLM for our test set, reported in LABEL:tab:llm_set, considering a fixed input length of $n_{\text{in}}=64$. The sub-figures correspond to different LLM size families (XS, S, M), and illustrate how Energy Efficiency (y-axis) varies with the output length $n_{\text{out}}$ (x-axis). As shown, different LLMs exhibit a substantial spread in $E_{\text{eff}}$. For instance, LLama 3.2 1B results the most efficient LLM, reaching 120.71 at $n_{\text{out}}=256$ and 51.90 at $n_{\text{out}}=4096$, whereas Falcon RW 7B results the least efficient, with a Energy Efficiency values of 18.78 and 2.18 at the same output lengths. While the absolute efficiency values differ across models due to architectural differences, the overall behavior is consistent: each model exhibits a non-linear curve with a distinct peak in efficiency.
@@
To highlight this common trend, we construct an aggregated heatmap (Figure 3) by applying min–max normalization to the efficiency values of each LLM and then averaging across the full set. The resulting 3D visualization reports the Energy Efficiency (z-axis), bounded between 0 and 1, and illustrates how it varies with different output lengths (x-axis) and input lengths (y-axis). This representation reveals a consistent non-linear pattern: the highest efficiency occurs at $n_{\text{in}}=64$ and $n_{\text{out}}\in[128,256]$, whereas the lowest efficient is observed for shortest output lengths $n_{\text{out}}=64$ combined with the longest input length $n_{\text{in}}=4096$. These results suggest the existence of “Sweet Spots” where LLMs achieve the best trade-off between energy consumption and token generated. Quantitatively, comparing for each model the worst efficient tokens combination, always located at the larger input and shorter output, with the highest efficiency Sweet Spot case (reported in Table 5), we observe on average a $33.41\times$ ($\pm 4.90$) increase in efficiency across models. In Table 5 we also report every $\theta$ for our Memory and FLOPs based model, and the estimated Sweet Spots, using the Equation 6.
@@
Model Baseline 1 Baseline 2 Baseline 3 Baseline 4 FLOPs based Memory and FLOPs based
@@
Figure 3. Aggregated Energy Efficiency Heatmap.
Heatmaps showing the model Energy Efficiency aggregated average.
7. Analysis and estimation of Energy Efficiency Sweet Spots
To estimate the Energy-per-Token $E_{\text{tok}}$ for each combination of $n_{\text{in}}$ and $n_{\text{out}}$, we fitted the state-of-the-art Baseline 1-4 and our proposed analytical models described in Table 1 with the results of the experiments of each LLM separately, using a non-linear least square method. In Table 3 we report the MAPE (Mean Absolute Percentage Error) for each analytical model fitted on the relative LLM.
From Table 3 we can notice that the proposed models FLOPs-based and Memory-and-FLOPs-based consistently achieves higher accuracy than the baselines for all the tested LLM inference models. In average, the Baseline 1 and 2 (Niu et al., 2025) achieves a MAPE of 130.45% and 130.49%, respectively. Baseline 3 (Wilhelm et al., 2025) improves the MAPE in average to the 106.25%, and Baseline 4 (Wilkins et al., 2024) further reduce it to 31.40%. This suggests that adding both $n_{\text{in}}$ and $n_{\text{out}}$ as fitting parameters improves the Energy efficiency estimate for LLM inference w.r.t. considering only $n_{\text{out}}$. The two proposed analytical models (FLOPS-based and Memory-and-FLOPs-based ) improves the MAPE over Baseline 4, reducing it to 2.27% $(\pm 0.61)$ and 1.79% $(\pm 0.61)$ respectively. This underlines that the proposed FLOPs based model derived from computational complexity estimates better captures the dependency of LLM inference Energy Efficiency w.r.t. SoA, and that this dependency is non linear with both quadratic and mixed terms dependencies. Furthermore, the proposed Memory-and FLOPs-based consistently reduces the MAPE over the FLOPs-based one with an average reduction of the 18.6%, underlining that memory access operations are important as FLOPs and the fifth term in the analytical model is important.
8. Current limitations and Future works
8.1. Experimental settings
@@
The experimental evaluation presented in this work is subject to some limitations that should be acknowledged. First, our analysis focuses on relatively small-scale models, which may not fully capture the performance characteristics of larger LLMs with substantially different parameter counts and memory footprints. Second, all experiments were conducted on a single hardware platform, limiting the generalizability of the results across heterogeneous systems with different compute, memory, and interconnect properties. Finally, we relied on a single serving framework, which constrains the scope of our conclusions, as alternative serving solutions may employ different optimization strategies and execution models that affect performance.
@@
Future work will address these limitations by extending the experimental study to a broader range of model sizes, including large-scale and emerging LLM architectures, and by evaluating performance across multiple hardware platforms. In addition, incorporating multiple serving frameworks will allow a more systematic comparison of software-level optimizations and requests scheduling strategies. Developing a more generalized performance model that jointly accounts for compute- and memory-bound costs would provide a more comprehensive framework for predicting LLM serving efficiency and for guiding optimization strategies across a wider range of models and deployment scenarios.
8.2. Batch size characterization
The proposed analytical model does not explicitly account for the batch size, instead assuming a fixed maximum batch size by setting max_batch_size to 1024 for all evaluations.
As shown in Figure 1, the number of concurrent requests has a strong impact on Energy Efficiency, particularly at peak values, as the engine is able to process a larger number of requests simultaneously.
@@ 8.2. Batch size characterization
Figure 4. Energy efficiency of Llama 3.2 1B varying maximum batch size considering different number of requests.
Energy efficiency of Llama 3.2 1B varying maximum batch size considering different number of requests.
To analyze the interaction between batch size and the number of requests, Figure 4 reports the Energy Efficiency of Llama 3.2 1B while varying the maximum batch size configuration (max_batch_size) across a range of values: 64, 128, 256, 512 and 1024. We tested these configurations for two different workloads, namely $N_{\text{req}}=128$ and $N_{\text{req}}=1000$, with the input sequence length fixed to $n_{\text{in}}=64$.
As shown in 4(a), setting a maximum batch size larger than the number of requests (128) does not provide any efficiency benefit. Conversely, smaller batch sizes, such as 64, lead to a significant reduction in efficiency, as they prevent the engine from forming larger batches, resulting in under utilization of computational resources.
When considering a higher number of requests ($N_{\text{req}}=1000$), as illustrated in 4(b), the maximum batch size has a strong effect on overall efficiency. In particular, the peak efficiency decreases from 120.71 with a batch size of 1024 to 48.90 when the batch size is reduced to 64. However, the impact of the batch size diminishes for longer output sequences. For instance, when $n_{\text{out}}=4096$, the efficiency gap between batch sizes of 256 and 1024 narrows to 47.27 and 51.91, respectively. This reduced sensitivity to batch size is attributable to the decode phase becoming dominant in requests with short inputs and long outputs, causing the engine to be constrained primarily by memory bandwidth rather than by the batch dimension.
In future work, we plan to extend our analysis of the impact of maximum batch size on LLM Energy Efficiency and to explicitly incorporate this parameter into the proposed analytical model.
@@ 9. Conclusions
This paper presented a comprehensive study of the Energy Efficiency of LLM inference evaluated on NVIDIA H100 GPUs with TensorRT-LLM. We showed that energy consumption during inference can be explained by decomposing FLOPs and Memory Access contributions into prefill and decode phases, and by normalizing them on a per-token basis. The resulting analytical model highlights the quadratic overhead of input processing, the amortization effect of output length, and the linear cost of sequential decoding. Our experiments confirm that this model aligns closely with real-world energy measurements. Specifically, we identified efficiency regimes where energy-per-token reaches a minimum, and demonstrated that these correspond to the balance point where input costs are sufficiently amortized without incurring excessive decoding overhead. This explains why efficiency peaks at short-to-moderate input lengths and medium outputs, while dropping significantly for long prompts or short generations.
Beyond theoretical insight, our findings have practical implications for LLM deployment. The analytical structure of inference costs allow to foresee efficiency Sweet Spots and to adapt prompt or generation length to reduce energy consumption. This perspective transforms Energy Efficiency from an empirical concept into a predictable design parameter. Building a generalized compute–memory efficiency model will further broaden the applicability, supporting sustainable optimization across different architectures and deployment contexts. In summary, our study represents an intermediate step between empirical benchmarking and theoretical modeling of LLM inference costs, providing a model that not only explains current efficiency regimes but also guides more energy-aware deployment strategies for large-scale generative AI.
@@ Appendix A Appendix
This appendix provides additional quantitative details supporting the modeling and statistical analysis presented in the main body of the paper. In particular, it reports the full set of regression parameters significance statistics and values for the proposed Memory and FLOPs based analytical model across all evaluated LLMs.
A.1. Parameters significance
Table 4 summarizes the aggregated statistical significance of each model parameter across all evaluated LLMs. By reporting the distribution of $p$-values for each parameter, this table highlights which components of the model consistently contribute to explaining performance variation and which parameters exhibit weaker or model-dependent influence. Together, these tables complement the main results by offering a detailed view of the model fitting behavior and its statistical robustness across different architectures and scales. In the table, the number asterisks quantify the significance of the parameter, with (***) meaning high significance, (**) moderate significance, (*) weaker significance and n.s. not significant.
Table 4. Aggregated statistical significance across LLM datasets.
Parameter $p<0.001$ (***) $p<0.01$ (**) $p<0.05$ (*) $p\geq 0.05$ (n.s.)
$\theta_{0}$ 13/13 (100%) 0/13 (0%) 0/13 (0%) 0/13 (0%)
$\theta_{1}$ 13/13 (100%) 0/13 (0%) 0/13 (0%) 0/13 (0%)
$\theta_{2}$ 13/13 (100%) 0/13 (0%) 0/13 (0%) 0/13 (0%)
$\theta_{3}$ 13/13 (100%) 0/13 (0%) 0/13 (0%) 0/13 (0%)
$\theta_{4}$ 13/13 (100%) 0/13 (0%) 0/13 (0%) 0/13 (0%)
$\theta_{5}$ 8/13 (62%) 1/13 (8%) 3/13 (23%) 1/13 (8%)
A.2. Parameters value
Table 5 lists the fitted values of the model parameters $\theta_{0}$$\theta_{5}$ for each LLM, grouped by scale. In addition, it reports both the empirically observed sweet spots and the corresponding values estimated by the proposed model, enabling a direct comparison between measured and predicted optimal configurations. These results provide transparency into the parameterization of the model and allow reproducibility of the sweet-spot estimation process.
Table 5. Theta values and Sweet-Spots estimation for our Memory and FLOPs based model. (Percentages may not sum to 100% due to rounding)
Model $\theta_{0}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ $\theta_{4}$ $\theta_{5}$ Sweet Spot Sweet Spot (Estimated)
XS models
Llama 3.2 (1B) 5.005153e-03 1.079941e-07 6.825240e-06 2.611042e-03 3.852659e-06 5.406443e-01 64/256 64/429
OPT (1.3B) 5.969229e-03 1.666292e-07 4.388860e-05 2.915735e-03 2.342971e-05 3.187084e-01 64/64 64/147
Qwen 2 (1.5B) 6.528509e-03 9.147483e-08 6.199954e-06 3.700520e-03 3.288754e-06 3.805096e-01 64/512 64/433
S models
Gemma 2 (2B) 9.468873e-03 1.223718e-07 2.486662e-05 5.371303e-03 1.372551e-05 5.856132e-01 64/256 64/260
OPT (2.7B) 6.541956e-03 2.855932e-07 9.197051e-05 5.909706e-03 5.038964e-05 8.566784e-01 64/128 64/157
Llama 3.2 (3B) 9.259155e-03 1.842970e-07 2.709305e-05 6.842019e-03 1.461407e-05 8.942031e-01 64/256 64/302
Granite (3B) 9.300161e-03 4.210179e-07 9.340093e-05 7.914036e-03 5.199433e-05 1.170166e+$0$ 64/128 64/180
M models
OPT (6.7B) 1.300082e-02 4.308046e-07 1.440099e-04 1.379572e-02 8.632801e-05 1.016691e+$0$ 64/128 64/148
Qwen 2 (7B) 1.857219e-02 2.598309e-07 1.531846e-05 1.465957e-02 9.735924e-06 8.719728e-01 64/512 64/431
Falcon-RW (7.5B) 1.871769e-02 9.200243e-07 1.652377e-04 1.673305e-02 1.052294e-04 7.410144e-01 64/128 64/131
Granite (8B) 2.073270e-02 3.926489e-07 4.029094e-05 1.744363e-02 2.512785e-05 8.540071e-01 64/256 64/280
Llama 3.1 (8B) 1.970934e-02 3.329326e-07 3.583505e-05 1.551942e-02 2.249613e-05 8.939092e-01 64/256 64/290
Gemma 2 (9B) 1.749322e-02 5.981696e-07 1.110917e-04 1.901597e-02 7.408055e-05 2.558409e+$0$ 64/64 64/226

### v2

@@ 5.2.3. Observation of Peak Energy Efficiency Spots
In Figure 3, we show the energy efficiency of each LLM in our test set, reported in LABEL:tab:llm_set, considering a fixed input length of $n_{\text{in}}=64$, the input length which shows the maximum efficiency, and varying the output length. The sub-figures correspond to different LLM size families (XS, S, M), and illustrate how energy efficiency (y-axis) varies with the output length $n_{\text{out}}$ (x-axis). We see that different LLMs exhibit a substantial spread in $E_{\text{eff}}$. For instance, LLama 3.2 1B is the most efficient LLM, reaching 120.71 at $n_{\text{out}}=256$ and 51.90 at $n_{\text{out}}=4096$, whereas Falcon-RW 7B is the least efficient, with energy efficiency values of 18.78 and 2.18 at the same output lengths. While the absolute efficiency values differ across models due to architectural differences, the overall behavior is consistent: each model exhibits a non-linear curve with a distinct peak in efficiency.
@@
To highlight this common trend, we construct an aggregated heatmap in Figure 4 by applying min–max normalization to the efficiency values of each LLM and then averaging across the full set. The resulting 3D visualization reports the energy efficiency (z-axis), bounded between 0 and 1, and illustrates how it varies with different output lengths (x-axis) and input lengths (y-axis). This representation reveals a consistent non-linear pattern: the highest efficiency occurs at $n_{\text{in}}=64$ and $n_{\text{out}}\in[128,256]$, whereas the lowest efficiency is observed for shortest output lengths $n_{\text{out}}=64$ combined with the longest input length $n_{\text{in}}=4096$. These results suggest the existence of “sweet spots” where LLMs achieve the best trade-off between energy consumption and tokens generated.
Quantitatively, comparing for each model the worst efficient tokens combination, always located at the larger input and shorter output, with the highest energy efficiency configuration case, we observe on average a $33.41\times$ ($\pm 4.90$) increase in efficiency across models. In Appendix A.2 (see Table 5), we report the $\theta$ parameters for our SweetSpot model, along with the experimental and the estimated peak efficiency configurations, calculated using Equation 6.
Figure 4. Aggregated energy efficiency heatmap.
Heatmaps showing the model Energy Efficiency aggregated average.
5.2.4. Prediction of Peak Energy Efficiency Spots
To estimate the Energy-per-Token $E_{\text{tok}}$ for each combination of $n_{\text{in}}$ and $n_{\text{out}}$, we fitted the state-of-the-art Baseline 1-4 models and our proposed analytical model SweetSpot described in Table 1 with the results of the experiments of each LLM separately, using a non-linear least square method. In Table 3, we report the Mean Absolute Percentage Error (MAPE) for each model fitted on the relative LLM.
We can observe that the proposed SweetSpot model consistently achieves higher accuracy than the baselines for all the tested LLMs. On average, the Baselines 1 and 2 (Niu et al., 2025) achieve a MAPE of 130.45% and 130.49%, respectively. Baseline 3 (Wilhelm et al., 2025) improves the MAPE on average to 106.25%, and Baseline 4 (Wilkins et al., 2024) further reduces it to 31.40%. This suggests that adding both $n_{\text{in}}$ and $n_{\text{out}}$ as fitting parameters improves the energy efficiency estimation of LLM inference w.r.t. considering only $n_{\text{out}}$. Our proposed model improves the MAPE over Baseline 4 significantly, reducing it to 2.27% $(\pm 0.61)$ and 1.79% $(\pm 0.61)$ in its FLOPs-only and complete versions respectively. This underlines that the FLOPs-only SweetSpot model derived from computational complexity estimates captures better the dependency of LLM inference energy efficiency w.r.t. SoA, and that this dependency is non-linear with both quadratic and mixed terms dependencies. Furthermore, the memory extended SweetSpot model consistently reduces the MAPE over the FLOPs-only one with an average reduction of 18.6%, underlining that memory-access operations are as important as FLOPs and the sixth term $\frac{\theta_{5}}{n_{\text{out}}}$ in the analytical model is crucial.
@@ 5.2.4. Prediction of Peak Energy Efficiency Spots
Model Baseline 1 Baseline 2 Baseline 3 Baseline 4 SweetSpot (FLOPs-only) SweetSpot
@@
6. Limitations and Future work
Although the SweetSpot model provides meaningful insights into LLM inference energy efficiency, its generality is limited by assumptions on model scale, hardware platform, inference framework, and batching configuration. This section discusses these limitations and outlines directions for future work.
6.1. Experimental Setting
@@
The current study focuses on not very large models, which may not fully capture the performance characteristics of larger LLMs with substantially different parameter counts and memory footprints. Second, all experiments were conducted on a single hardware platform, limiting the generalizability of the results to heterogeneous systems with different compute, memory, and interconnect properties. Finally, we relied on a single inference framework, which constrains the scope of our conclusions, as alternative inference solutions may employ different optimization strategies and execution models affecting performance.
@@
Future work will address these limitations by extending the experimental study to a broader range of model sizes, including large-scale and emerging LLM architectures, and by evaluating performance across multiple hardware platforms. In addition, incorporating multiple inference frameworks will allow a more systematic comparison of software-level optimizations and request scheduling strategies. Developing such a generalized performance model would provide a more comprehensive framework for predicting LLM inference efficiency and for guiding optimization strategies across a wider range of models and deployment scenarios.
6.2. Batch Size Impact
The proposed SweetSpot analytical model does not explicitly account for the batch size, instead assumes a fixed maximum batch size, which is set to 1024 in all evaluations. As shown in Figure 1, the number of concurrent requests has a strong impact on energy efficiency, particularly at peak values, as the engine can process a larger number of requests simultaneously. To analyze the interaction between the batch size and the number of requests, in Figure 5 we report the energy efficiency of Llama 3.2 1B (the most efficient LLM in the test set) while varying the maximum batch size configuration (max_batch_size) across a range of values: 64, 128, 256, 512 and 1024. We tested these configurations for $N_{\text{req}}=128$ and $N_{\text{req}}=1000$, with the input sequence length fixed to $n_{\text{in}}=64$, where the maximum peak of energy efficiency occurs.
@@ 6.2. Batch Size Impact
Figure 5. Energy efficiency of Llama 3.2 1B varying maximum batch size considering different numbers of requests.
[…]
In Figure 5a, we observe that setting a maximum batch size larger than $N_{\text{req}}=128$ does not provide any efficiency benefit. Conversely, a smaller batch size like 64 leads to a significant reduction in efficiency, as this prevents the engine from forming larger batches, resulting in under utilization of computational resources.
When considering a higher number of requests ($N_{\text{req}}=1000$), as illustrated in Figure 5b, the maximum batch size has a strong effect on overall efficiency. In particular, the peak efficiency decreases from 120.71 with a batch size of 1024 to 48.90 when the batch size is reduced to 64. However, the impact of the batch size diminishes for longer output sequences. For instance, when $n_{\text{out}}=4096$, the efficiency gap between batch sizes of 256 and 1024 narrows to 47.27 and 51.91, respectively. This reduced sensitivity to batch size is attributable to the decode phase becoming dominant in requests with short inputs and long outputs, causing the engine to be constrained primarily by memory bandwidth rather than by the batch dimension.
In future work, we will extend our analysis of how maximum batch size affects LLM energy efficiency and explicitly incorporate this parameter into the proposed analytical model.
@@ 7. Conclusions
This paper presented a comprehensive study of the energy efficiency of LLM inference evaluated on NVIDIA H100 GPUs with TensorRT-LLM. We showed that energy consumption during inference can be explained by decomposing FLOPs and memory-access contributions into prefill and decode phases, and by normalizing them on a per-token basis. The resulting analytical model, SweetSpot, highlights the quadratic overhead of input processing, the amortization effect of output length, and the linear cost of sequential decoding. Our experiments confirm that SweetSpot aligns closely with real-world energy measurements. Specifically, we identified efficiency regimes where energy-per-token reaches a minimum, and demonstrated that these correspond to the balance point where input costs are sufficiently amortized without incurring excessive decoding overhead. This explains why efficiency peaks at short-to-moderate input lengths and medium outputs, while dropping significantly for long prompts or short generations.
Beyond theoretical insight, our findings have practical implications for LLM deployment. The analytical structure of inference costs allow to foresee energy efficiency ”sweet-spots” configurations and to adapt prompt or generation length to reduce energy consumption. This perspective transforms energy efficiency from an empirical concept into a predictable design parameter. Building a generalized compute–memory efficiency model will further broaden the applicability, supporting sustainable optimization across different architectures and deployment contexts. In summary, our study represents an intermediate step between empirical benchmarking and theoretical modeling of LLM inference costs, providing a model that not only explains current efficiency regimes but also guides more energy-aware deployment strategies for large-scale generative AI.
Acknowledgements.
The research has been partially funded by the EU Pilot for exascale EuroHPC EUPEX (g.a. 101033975), EuroHPC JU SEANERGYS (g.a. 101177590), DARE (g.a. 101143421) and Spoke “FutureHPC & BigData” of the ICSC-Centro Nazionale di Ricerca in “High Performance Computing, Big Data & Quantum Computing”, funded by the EU - NextGenerationEU projects.
@@ Appendix A Appendix
This appendix provides additional quantitative details supporting the modeling and statistical analysis presented in the main body of the paper. In particular, it reports the complete set of regression parameter significance statistics and values for our proposed SweetSpot analytical model across all evaluated LLMs.
A.1. Parameter Significance
Table 4 summarizes the aggregated statistical significance of each analytical model parameter across all evaluated LLMs. By reporting the distribution of $p$-values for each parameter, this table highlights which components of the model consistently contribute to explaining performance variation and which parameters exhibit weaker or model-dependent influence. Together, these tables complement the main results by offering a detailed view of the model fitting behavior and its statistical robustness across different architectures and scales. In the table, the number of asterisks quantify the significance of the parameter, with (***) meaning high significance, (**) moderate significance, (*) weaker significance and (n.s.) not significant.
A.2. Parameter Values
Table 5 lists the fitted values of the model parameters $\theta_{0}$$\theta_{5}$ for each LLM, grouped by scale. In addition, it reports both the empirically observed peak energy-efficiency configurations $(n_{\text{in}},n_{\text{out}})$ and the corresponding values estimated by our SweetSpot model, enabling a direct comparison between measured and predicted optimal configurations. These results provide transparency into the parameterization of the model and allow reproducibility of the SweetSpot prediction process.